At 7 feet 8 inches, the world's tallest woman casts a 46-inch shadow. At the same time, the world's shortest woman casts a 15.5 inch shadow. How tall is the world's shortest woman?
Answer by TimothyLamb(4379)

(7*12+8)/46 = x/15.5
x = 15.5*(7*12+8)/46
x = 31 inches
x = 2.58333333333 feet
x = 2 feet (.58333333333*12) inches
---
answer:
x = 2 feet 7 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php